# AlignBot: Aligning VLM-powered Customized Task Planning with User Reminders Through Fine-Tuning for Household Robots

### v1

@@ I Introduction
Vision-Language Models (VLMs) are emerging as a new paradigm in robotic task planning [5, 6, 7, 8, 9, 10, 11, 12]. Compared to Large Language Models (LLMs), VLMs can leverage both visual and language information, and obtain a more comprehensive understanding of the task context. However, aligning VLM-powered task planners for household robots with user reminders is challenging due to the limited quantity of reminders, their diversity in type, and their multimodal nature. In a home setting, user-robot interactions are relatively infrequent, resulting in a sparse reminder dataset, which makes it extremely challenging to fine-tune a VLM with a large number of parameters, such as GPT-4o [13]. Moreover, user reminders are inherently diverse, including personalized preferences, corrective guidance, and contextual assistance [14, 15, 2, 16, 3, 17], as illustrated in Fig. 1. This diversity necessitates a model with strong generalization capabilities to effectively process and learn from the varied reminders. Additionally, the multimodal nature of user reminders, which frequently involves both text and visual elements that are closely linked, increases the alignment complexity, posing challenges that single-modality models cannot fully address.
@@
To this end, we introduce AlignBot, an advanced framework designed to effectively align diverse and multimodal user reminders with robotics customized task planning for everyday household activities. Within AlignBot, a fine-tuned LLaVA-7B model functions as an adapter for GPT-4o [25, 13, 26, 27], internalizing user reminders to generate instruction-formatted cues that guide GPT-4o in task planning. By “freezing” GPT-4o and fine-tuning only the much smaller LLaVA-7B model, we significantly reduce data requirements while leveraging LLaVA-7B’s multimodal capabilities to handle multimodal reminders. Within our data collection pipeline, we address the diversity of user-robot interactions by generating standardized instruction-formatted cues that transform complex dialogue inputs into consistent prompts. These cues streamline the processing of diverse reminders and optimize GPT-4o’s performance, given its strength in interpreting and executing structured prompts [28]. To address action omissions and sequence errors in long-term tasks, we also implement a dynamic retrieval mechanism that selects relevant cases from historically successful task plans as prompts for GPT-4o, thereby improving task execution accuracy.
@@ IV-A Fine-Tuning LLaVA with User Reminders
LLaVA Fine-Tuning: For fine-tuning, we employ the LLaVA-7B model [^3] https://huggingface.co/llava-hf/llava-1.5-7b-hf, integrating it with GPT-4o by freezing the GPT-4o parameters and exclusively fine-tuning LLaVA. This approach preserves GPT’s task planning capabilities while enabling LLaVA to achieve better performance in both semantic grounding and cue generation tasks. LLaVA-7B is selected over LLaVA-13B due to its lower data requirements, rendering it more suitable for contexts with limited training data. Although LLaVA-7B exhibits comparatively weaker reasoning capabilities than LLaVA-13B, the latter demonstrates no significant performance gains when trained on small datasets, thereby making LLaVA-7B a more computationally efficient choice under our constraints. To further optimize LLaVA-7B’s fine-tuning, we leverage the LoRA method [26]. This method enables efficient fine-tuning by reducing the number of trainable parameters, allowing the model to swiftly capture critical details even in low-data settings.
@@
The fine-tuning process is optimized with the Cross-Entropy Loss function [26]. The number of epochs is carefully set to 40 for our fine-tuning process, representing a substantial increase compared to the majority of existing studies, which typically employ fewer than 10 epochs. This decision is motivated by the specific context of our work: a single household environment with task repetition and user-specific details. Although increasing the number of epochs can reduce generalization, this trade-off is acceptable in our case, as the model needed to retain detailed information for long-term adaptation to a particular user. We also explore the cold-start problem, particularly when working with very small datasets (e.g., dozens of samples). In such cases, we increase the number of epochs to 60 to help the model memorize the initial reminder. As the dataset size expands, we gradually reduce the epochs back to 40 to ensure the model can adapt to new tasks without losing the ability to recall past interactions. We experimented with learning rates ranging from $1\times 10^{-5}$ to $5\times 10^{-5}$ and found no significant differences in performance across this range. Further details on hyperparameters are provided in Appendix B.
@@ IV-B Case-Based Learning for Enhanced GPT Prompting
Long-horizon task planning in household robotics frequently encounters issues such as action omissions and sequence errors, undermining execution reliability. For example, a robot may attempt to place an object without first executing the necessary action of grasping it. To address these challenges, we propose a case-based learning approach that enhances GPT-4o’s planning capabilities by incorporating relevant historical cases into the prompting process. This method involves dynamically retrieving the top-$k$ most relevant cases-specifically, past action plans that were both successful and approved by users-from historical dialogue data. Here, we refer to these as cases. The retrieved cases serve as reference examples for GPT-4o, providing contextual guidance based on prior successful executions. By leveraging similar and correct action sequences from past cases, we aim to improve the accuracy and robustness of the generated plans. The rationale behind this approach is that previous successful plans encapsulate effective solutions that can help prevent common errors in new tasks. However, retrieving relevant cases in a multimodal context poses challenges due to the complexity of assessing similarity across different modalities [41]. To overcome this, we first select successful plans that match the same task description. We then utilize GPT-4o’s feedback to evaluate and rank these plans based on their relevance and effectiveness. By dynamically integrating the most appropriate cases as references, GPT-4o can generate more accurate and comprehensive action sequences, thereby reducing omissions and enhancing overall task execution.
@@
Consider a subset $D_{r}$ of raw dataset $D$, defined as $D_{r}=\{\langle p_{1},t_{1}\rangle,\langle p_{2},t_{2}\rangle,\dots\}$, where each action plan $p_{i}$ is linked to a historical corresponding task description $t_{i}$. Each plan $p_{i}$ is assigned an initial priority score $f_{i}=f_{0}$. When a new task $t_{\text{new}}$ is introduced, the system selects a subset $D_{r}^{\prime}\subset D_{r}$ of action plans relevant to $t_{\text{new}}$, determined by the similarity function $sim(t_{i},t_{\text{new}})$. Specifically, $D_{r}^{\prime}=\{p_{i}\mid sim(t_{i},t_{\text{new}})>\tau\}$, where $\tau$ is a predefined similarity threshold. Here, the similarity function is computed using TF-IDF [42] combined with cosine similarity. The action plans in $D_{r}^{\prime}$ are then sorted based on their priority scores $f_{i}$, resulting in an ordered set. To normalize these priority scores and prevent over-reliance on a limited set of action plans, a Softmax function is applied: $f_{i_{j}}=\exp(\hat{f}_{i_{j}})/\sum_{k=1}^{|S|}\exp(\hat{f}_{i_{k}})$. To dynamically adjust priorities based on effectiveness, we introduce a gradient-based update mechanism. Each action plan $p_{i}$ is associated with a gradient value $\Delta f$, representing the rate of change in its priority. The gradient is updated depending on the utility of the action plan in the current task: $\Delta f_{\text{positive}}=\Delta f_{\text{positive}}^{0}\cdot\exp(-\alpha\cdot n_{i_{j}})$, $\Delta f_{\text{negative}}=\Delta f_{\text{negative}}^{0}\cdot\exp(-\beta\cdot n_{i_{j}})$, where $\alpha$ and $\beta$ are decay factors controlling how quickly priorities change with repeated use, and $n_{i_{j}}$ denotes the usage frequency of plan $p_{i}$. The initial gradient $\Delta f^{0}$ is a fixed value controlling the amplitude of updates, preventing unbounded growth or decay of priority scores over time. After updating the gradient, the priority score $f_{i}$ is adjusted accordingly: $f_{i}=\min(\max(f_{i}+\Delta f,0.1),1.0)$. The selection of action plans is governed by an $\epsilon$-Greedy strategy to balance exploration and exploitation. Specifically, with probability $\epsilon$, a random action plan $F_{\text{random}}$ is chosen to encourage exploration, while with probability $1-\epsilon$, the highest-priority plan $F_{\text{max}}$ is selected to exploit known effective strategies. The parameter $\epsilon$ controls this balance, ensuring adaptability without over-reliance on a limited set of strategies. By iteratively applying this method, the system refines its task planning, continuously adapting to user preferences and specific task contexts. Further details can be found in Appendix B.

### v2

@@ I Introduction
Vision-Language Models (VLMs) are emerging as a new paradigm in robotic task planning [5, 6, 7, 8, 9, 10, 11, 12, 13]. Compared to Large Language Models (LLMs), VLMs can leverage both visual and language information, and obtain a more comprehensive understanding of the task context. However, aligning VLM-powered task planners for household robots with user reminders is challenging due to the limited quantity of reminders, their diversity in type, and their multimodal nature. In a home setting, user-robot interactions are relatively infrequent, resulting in a sparse reminder dataset, which makes it extremely challenging to fine-tune a VLM with a large number of parameters, such as GPT-4o [14]. Moreover, user reminders are inherently diverse, including personalized preferences, corrective guidance, and contextual assistance [15, 16, 2, 17, 3, 18], as illustrated in Fig. 1. This diversity necessitates a model with strong generalization capabilities to effectively process and learn from the varied reminders. Additionally, the multimodal nature of user reminders, which frequently involves both text and visual elements that are closely linked, increases the alignment complexity, posing challenges that single-modality models cannot fully address.
@@
To this end, we introduce AlignBot, an advanced framework designed to effectively align diverse and multimodal user reminders with robotics customized task planning for everyday household activities. Within AlignBot, a fine-tuned LLaVA-7B model functions as an adapter for GPT-4o [26, 14, 27, 28, 29, 30], internalizing user reminders to generate instruction-formatted cues that guide GPT-4o in task planning. By “freezing” GPT-4o and fine-tuning only the much smaller LLaVA-7B model, we significantly reduce data requirements while leveraging LLaVA-7B’s multimodal capabilities to handle multimodal reminders. Within our data collection pipeline, we address the diversity of user-robot interactions by generating standardized instruction-formatted cues that transform complex dialogue inputs into consistent prompts. These cues streamline the processing of diverse reminders and optimize GPT-4o’s performance, given its strength in interpreting and executing structured prompts [31]. To address action omissions and sequence errors in long-term tasks, we also implement a dynamic retrieval mechanism that selects relevant cases from historically successful task plans as prompts for GPT-4o, thereby improving task execution accuracy.
@@ IV-A Fine-Tuning LLaVA with User Reminders
LLaVA Fine-Tuning: For fine-tuning, we employ the LLaVA-7B model [^3] https://huggingface.co/llava-hf/llava-1.5-7b-hf, integrating it with GPT-4o by freezing the GPT-4o parameters and exclusively fine-tuning LLaVA. This approach preserves GPT’s task planning capabilities while enabling LLaVA to achieve better performance in both semantic grounding and cue generation tasks. LLaVA-7B is selected over LLaVA-13B due to its lower data requirements, rendering it more suitable for contexts with limited training data. Although LLaVA-7B exhibits comparatively weaker reasoning capabilities than LLaVA-13B, the latter demonstrates no significant performance gains when trained on small datasets, thereby making LLaVA-7B a more computationally efficient choice under our constraints. To further optimize LLaVA-7B’s fine-tuning, we leverage the LoRA method [27]. This enables efficient fine-tuning by reducing the number of trainable parameters, allowing the model to swiftly capture critical details even in low-data settings.
@@
The fine-tuning process is optimized with the Cross-Entropy Loss function [27, 44]. The number of epochs is carefully set to 40 for our fine-tuning process, representing a substantial increase compared to the majority of existing studies, which typically employ fewer than 10 epochs. This decision is motivated by the specific context of our work: a single household environment with task repetition and user-specific details. Although increasing the number of epochs can reduce generalization, this trade-off is acceptable in our case, as the model needed to retain detailed information for long-term adaptation to a particular user. We also explore the cold-start problem, particularly when working with very small datasets (e.g., dozens of samples). In such cases, we increase the number of epochs to 60 to help the model memorize the initial reminder. As the dataset size expands, we gradually reduce the epochs back to 40 to ensure the model can adapt to new tasks without losing the ability to recall past interactions. We experimented with learning rates ranging from $1\times 10^{-5}$ to $5\times 10^{-5}$ and found no significant differences in performance across this range. Further details on hyperparameters are provided in Appendix B.
@@ IV-B Case-Based Learning for Enhanced GPT Prompting
Long-horizon task planning in household robotics frequently encounters issues such as action omissions and sequence errors, undermining execution reliability. For example, a robot may attempt to place an object without first executing the necessary action of grasping it. To address these challenges, we propose a case-based learning approach that enhances GPT-4o’s planning capabilities by incorporating relevant historical cases into the prompting process. This method involves dynamically retrieving the top-$k$ most relevant cases-specifically, past action plans that were both successful and approved by users-from historical dialogue data. Here, we refer to these as cases. These cases provide GPT-4o with contextual guidance from past successes. By using similar, correct past action sequences, we aim to improve plan accuracy and robustness. The rationale is that past successful plans contain effective solutions to prevent common errors. However, retrieving relevant cases in a multimodal context poses challenges due to the complexity of assessing similarity across different modalities [45]. To overcome this, we first select successful plans that match the same task description. We then utilize GPT-4o’s feedback to evaluate and rank these plans based on their relevance and effectiveness. By integrating the most appropriate cases, GPT-4o generates more accurate and comprehensive action sequences, reducing omissions and improving execution.
@@
Consider a subset $D_{r}$ of raw dataset $D$, defined as $D_{r}=\{\langle p_{1},t_{1}\rangle,\langle p_{2},t_{2}\rangle,\dots\}$, where each action plan $p_{i}$ is linked to a historical corresponding task description $t_{i}$. Each plan $p_{i}$ is assigned an initial priority score $f_{i}=f_{0}$. When a new task $t_{\text{new}}$ is introduced, the system selects a subset $D_{r}^{\prime}\subset D_{r}$ of action plans relevant to $t_{\text{new}}$, determined by the similarity function $sim(t_{i},t_{\text{new}})$. Specifically, $D_{r}^{\prime}=\{p_{i}\mid sim(t_{i},t_{\text{new}})>\tau\}$, where $\tau$ is a predefined similarity threshold. The similarity is computed using TF-IDF [46] and cosine similarity. The action plans in $D_{r}^{\prime}$ are then sorted based on their priority scores $f_{i}$, resulting in an ordered set. To normalize these priority scores and prevent over-reliance on a limited set of action plans, a Softmax function is applied: $f_{i_{j}}=\exp(\hat{f}_{i_{j}})/\sum_{k=1}^{|S|}\exp(\hat{f}_{i_{k}})$. To dynamically adjust priorities based on effectiveness, we introduce a gradient-based update mechanism. Each plan $p_{i}$ has a gradient $\Delta f$, indicating its priority change rate. The gradient is updated depending on the utility of the action plan in the current task: $\Delta f_{\text{positive}}=\Delta f_{\text{positive}}^{0}\cdot\exp(-\alpha\cdot n_{i_{j}})$, $\Delta f_{\text{negative}}=\Delta f_{\text{negative}}^{0}\cdot\exp(-\beta\cdot n_{i_{j}})$, where $\alpha$ and $\beta$ are decay factors controlling how quickly priorities change with repeated use, and $n_{i_{j}}$ denotes the usage frequency of plan $p_{i}$. The initial gradient $\Delta f^{0}$ is a fixed value controlling the amplitude of updates, preventing unbounded growth or decay of priority scores over time. After updating the gradient, the priority score $f_{i}$ is adjusted accordingly: $f_{i}=\min(\max(f_{i}+\Delta f,0.1),1.0)$. The selection of action plans is governed by an $\epsilon$-Greedy strategy to balance exploration and exploitation. Specifically, with probability $\epsilon$, a random action plan $F_{\text{random}}$ is chosen to encourage exploration, while with probability $1-\epsilon$, the highest-priority plan $F_{\text{max}}$ is selected to exploit known effective strategies. The parameter $\epsilon$ controls this balance, ensuring adaptability without over-reliance on a limited set of strategies. By iteratively applying this method, the system refines its task planning, continuously adapting to user preferences and specific task contexts. Further details can be found in Appendix B.
@@ VI Conclusion
Acknowledgment
This work is supported by the Shanghai AI Laboratory, the National Key R&D Program of China (2022ZD0160102), the National Natural Science Foundation of China (62376222), and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).